OpenStudy (dunk789):

let f(x)=x-3 and g(x)=2f(x). then g(2=

OpenStudy (will.h):

try to enhance typing the question it just doesn't make sense right now

OpenStudy (dunk789):

let f(x)=x-3 and g(x)=2f(x). then g(2)=______

OpenStudy (liziekay):

if f(x)= x-3 then g(x)= 2f(x). so when they ask for g(2) you just substitute for x, g(2)= 2(2-3) => g(2) = -2

OpenStudy (will.h):

great! So the function g(x) is 2 times f(x) so g(x) would be 2x - 6 and to find g(2) plug in x = 2 in g(x) function that we just made and see the value. g(2) = -2

OpenStudy (dunk789):

ohh i see

OpenStudy (will.h):

If you need help with other question let me know :)